# See2Refine: Vision-Language Feedback Improves LLM-Based eHMI Action Designers

### v1

@@ A.4.3 Lightbar Modality Performance
For the lightbar modality condition with 9 participants, repeated measures ANOVA also revealed significant main effects for five of six metrics:
Message Targeting. The ANOVA showed a significant main effect ($F(4,32)=6.34$, $p<0.001$, $\eta^{2}=0.442$). Grok 4 ($M=5.35$, $SD=1.03$), OpenAI GPT-5 ($M=5.31$, $SD=1.08$), and Claude Opus 4.1 ($M=5.32$, $SD=0.95$) achieved the highest ratings. DesignerLLM ($M=4.75$, $SD=1.39$) demonstrated a 12.6% improvement over the base model ($M=4.22$, $SD=1.33$).
@@
User Acceptance. The ANOVA indicated highly significant effects ($F(4,32)=13.55$, $p<0.001$, $\eta^{2}=0.629$), representing the largest effect size across all metrics and modalities. OpenAI GPT-5 ($M=5.31$, $SD=1.47$), Claude Opus 4.1 ($M=5.29$, $SD=1.29$), and Grok 4 ($M=5.06$, $SD=1.58$) achieved the highest acceptance ratings. DesignerLLM ($M=4.33$, $SD=1.66$) achieved a substantial 12.5% improvement over the base model ($M=3.85$, $SD=1.75$), representing the largest single-metric gain.
Consistency. Significant model differences were found ($F(4,32)=7.54$, $p<0.001$, $\eta^{2}=0.485$). OpenAI GPT-5 ($M=5.40$, $SD=1.39$) and Claude Opus 4.1 ($M=5.28$, $SD=1.31$) achieved the highest ratings. DesignerLLM ($M=4.39$, $SD=1.23$) demonstrated 3.5% improved consistency over the base model ($M=4.24$, $SD=1.15$).
@@ A.4.4 Summary
The human evaluation results demonstrate that our co-learning framework successfully enhances a 7B parameter model (DesignerLLM) to generate significantly more acceptable eHMI actions compared to its base model counterpart (Qwen2.5-7B-Instruct). Across both modalities, DesignerLLM achieved an average improvement of 7.8% across all metrics, with the most substantial gains observed in message targeting (15.4% for eye, 12.6% for lightbar) and user acceptance (5.0% for eye, 12.5% for lightbar). While commercial models with substantially larger parameter counts (GPT-5, Grok 4, Claude Opus 4.1) achieved higher absolute ratings, DesignerLLM’s consistent rank-4 performance across both modalities and all six metrics validates the effectiveness of our framework. Notably, the large effect sizes observed across most metrics ($\eta^{2}$ ranging from 0.318 to 0.629) indicate that model choice has a substantial impact on perceived eHMI quality. These results confirm that our framework not only improves automated VLM-based evaluation scores but also translates to meaningful enhancements in real human user experience, achieving competitive performance with a fraction of the computational resources required by commercial alternatives.
Figure 7: Boxplot of metric distributions at different training milestones. User Acceptance and Consistency both show increasing averages, with their distributions becoming more concentrated. In contrast, Trust exhibits minimal change.

### v2

@@ A.4.3 Lightbar Modality Performance
For the lightbar modality condition with nine participants, repeated–measures ANOVA also revealed significant main effects for five of six metrics:
Message Targeting. ANOVA showed a significant main effect ($F(4,32)=6.34$, $p<0.001$, $\eta^{2}=0.442$). Grok 4 ($M=5.35$, $SD=1.03$), OpenAI GPT-5 ($M=5.31$, $SD=1.08$), and Claude Opus 4.1 ($M=5.32$, $SD=0.95$) achieved the highest ratings. DesignerLLM ($M=4.75$, $SD=1.39$) demonstrated a 12.6% improvement over the base model ($M=4.22$, $SD=1.33$).
@@
User Acceptance. ANOVA indicated a highly significant effect ($F(4,32)=13.55$, $p<0.001$, $\eta^{2}=0.629$), representing the largest effect size across all metrics and modalities. OpenAI GPT-5 ($M=5.31$, $SD=1.47$), Claude Opus 4.1 ($M=5.29$, $SD=1.29$), and Grok 4 ($M=5.06$, $SD=1.58$) achieved the highest acceptance ratings. DesignerLLM ($M=4.33$, $SD=1.66$) achieved a substantial 12.5% improvement over the base model ($M=3.85$, $SD=1.75$), representing the largest single-metric gain.
Consistency. Significant model differences were found ($F(4,32)=7.54$, $p<0.001$, $\eta^{2}=0.485$). OpenAI GPT-5 ($M=5.40$, $SD=1.39$) and Claude Opus 4.1 ($M=5.28$, $SD=1.31$) achieved the highest ratings. DesignerLLM ($M=4.39$, $SD=1.23$) demonstrated a 3.5% improvement in consistency over the base model ($M=4.24$, $SD=1.15$).
@@ A.4.4 Summary
The human evaluation results demonstrate that our co-learning framework successfully enhances a 7B parameter model (DesignerLLM) to generate significantly more acceptable eHMI actions compared to its base model counterpart (Qwen2.5-7B-Instruct). Across both modalities, DesignerLLM achieved an average improvement of 7.8% across all metrics, with the largest gains observed in message targeting (15.4% for the eye and 12.6% for the lightbar) and user acceptance (5.0% for the eye and 12.5% for the lightbar). Although commercial models with substantially larger parameter counts, such as GPT-5, Grok 4, and Claude Opus 4.1, achieved higher absolute ratings, DesignerLLM consistently ranked fourth across both modalities and all six metrics, validating the effectiveness of our framework. Notably, the large effect sizes observed across most metrics ($\eta^{2}$ ranging from 0.318 to 0.629) indicate that model choice has a substantial impact on perceived eHMI quality. These results confirm that our framework not only improves automated VLM-based evaluation scores but also translates into meaningful enhancements in real human user experience, achieving competitive performance with a fraction of the computational resources required by commercial alternatives.
Appendix B Evaluator Prompts and Templates
This appendix provides the full prompts used in the multi-metric evaluation system (Section 3.3). All evaluations use GPT-5-mini OpenAI (2025). The video input consists of up to 50 frames, each resized to $512\!\times\!512$ pixels.
B.1 Phase 1: Pre-Evaluation Prompt (Without Intended Message)
B.1.1 System prompt
You are a communication receiver in a traffic condition. The detailed situation you encounter is as follows: {receiver_scenario} You are conducting an experiment to evaluate the message conveyance capability of the eHMI action (movement). You can clearly notice there is an external Human-Machine Interface (eHMI) installed on this autonomous vehicle. The eHMI type you see installed on the car is: {ehmi_short_description} What You Will Receive: Video frames showing eHMI behavior and animations. The video is rendered in Blender, but please treat it as a real-life scenario. The video will provide correct visual effects, lighting, and viewing angle from your perspective. The background is removed for fast rendering, so please ignore background effects and context when making your decision, focusing only on the action (movement) of the eHMI. Specific evaluation questionnaires to guide your assessment. What You Will Not Receive: The intended message that the vehicle equipped with the specific eHMI intends to convey to you. You are completing this questionnaire without knowing the intended message to ensure an unbiased evaluation based purely on your natural interpretation of the eHMI behavior. Your Role and Approach. You should: Act as a real human receiver in a traffic condition, considering factors that might affect human raters, including but not limited to: visual attractiveness, clarity, timing, message effectiveness, friendliness, and safety perception. Fully immerse yourself in the provided scenario, thinking about what a human communication receiver would think and need in this situation. Provide your immediate reaction as you would naturally respond in real traffic conditions. Give detailed reasoning for your ratings, explaining what specific elements influenced your assessment. Consider real-world implications of how this eHMI communication might affect traffic safety and human behavior.
B.1.2 User prompt
Based on the scenario and provided eHMI action (movement) video, please analyze what message you believe the eHMI is trying to convey, assess whether this communication is directed at you, and evaluate your trust in this autonomous vehicle system. Intention Recognition. Please carefully observe the eHMI behavior and infer what message the autonomous vehicle is attempting to communicate to you as a traffic participant. Consider the following aspects: visual patterns, movements, and animations; timing and duration of the eHMI behavior; color changes, brightness, or other visual cues; the context of the traffic scenario; what action or understanding the vehicle might be seeking from you; vehicle’s relative position and movement direction in relation to your location; and whether the eHMI signals appear to be directed specifically at you or other traffic participants. Rate your certainty in your message interpretation on a 9-point scale: 1 = completely uncertain $\to$ 5 = neutral certainty $\to$ 9 = completely certain. Message Targeting Assessment. Please assess whether you believe the eHMI communication is specifically directed at you as a traffic participant. Rate your confidence on a 9-point scale: 1 = definitely not for me $\to$ 5 = uncertain targeting $\to$ 9 = definitely for me. Trust Evaluation. Please evaluate the degree to which you believe the autonomous vehicle and its eHMI are reliable and will act safely as indicated. Rate your trust on a 9-point scale: 1 = complete distrust $\to$ 5 = neutral trust $\to$ 9 = complete trust.
B.2 Phase 2: Post-Evaluation Prompt (With Intended Message)
B.2.1 System prompt.
The setup is identical to Phase 1 (Appendix B.1), except the clause “What You Will Not Receive” is replaced with “What You Will Receive”.
B.2.2 User prompt.
The intended message is as follows: {intended_message} You are provided with this message after you experience this scenario (watch the video). Based on the scenario, provided eHMI action (movement) video, and the intended message the eHMI action wants to convey, please answer the following questions: User Acceptance Evaluation. Please evaluate the degree to which you would accept and approve of this eHMI-equipped autonomous vehicle based on the behavior shown in the video frames. Rate the overall user acceptance on a 9-point scale: 1 = strongly reject $\to$ 5 = mixed acceptance $\to$ 9 = fully accept. Consistency Evaluation. Please evaluate the consistency between the eHMI actions shown in the video frames and the intended message described above. Rate the overall consistency on a 9-point scale: 1 = completely inconsistent $\to$ 5 = mixed consistency $\to$ 9 = perfectly consistent.
B.3 Similarity Judge Prompt
After Phase 1, the VLM-interpreted message is compared with the ground-truth intended message using a small LLM judge (GPT-5-mini).
You are evaluating the similarity between an intended message and an interpreted message from an eHMI (external Human-Machine Interface) interaction. Task: Compare the intended message (ground truth) with the interpreted message (what the VLM guessed) and provide a similarity score. Scoring Criteria. Focus on the main meaning and urgency level: 1. Main Action Similarity: Do both messages convey the same core action/intention? (e.g., yielding, stopping, proceeding, waiting, warning) 2. Urgency Level: Do both messages show similar urgency or immediacy? 3. Overall Meaning Alignment: Do the messages fundamentally communicate the same thing to a traffic participant? Input: Intended Message: {intended_message} Interpreted Message: {interpreted_message}
Appendix C Training Hyperparameters
All fine-tuning experiments use LLaMA-Factory (Zheng et al., 2024) with LoRA (Hu et al., 2022) in bfloat16 precision. The data are formatted in ShareGPT style (Chiang et al., 2023) and split into training and test sets at an 80/20 ratio. Table 3 lists the full configuration for both stages.
Table 3: Training hyperparameters for Format-aware Fine-tuning (SFT) and Iterative Preference-based Learning (DPO).
Parameter SFT DPO
Base models Qwen2.5-7B / 1.5B-Instruct
Fine-tuning method LoRA LoRA
LoRA rank 8 8
Learning rate $5.0\!\times\!10^{-5}$ $5.0\!\times\!10^{-6}$
LR scheduler cosine cosine
Warmup ratio 0.1 0.1
Epochs 3 1
Effective batch size 16 16
Cutoff length 2048 2048
Precision bf16 bf16
DPO-specific
   DPO $\beta$ 0.2
   DPO loss sigmoid
   SFT mixing weight 0.01
Preference pair construction.
For each scenario, pairs $(y^{+},y^{-})$ are formed using the kernel score $K$. We first select the max–min pair when $\Delta K_{i}\geq\delta_{\min}$, where the default $\delta_{\min}=4.0$. We then rank the remaining high-gap pairs within the same scenario by score difference and retain the top 30%.
Quantization for local inference.
During DPO candidate generation, models are loaded with 4-bit NormalFloat (NF4) quantization.
Figure 7: Boxplots of User Acceptance, Consistency, and Trust across training milestones. User Acceptance and Consistency increase over time and show reduced variability, while Trust remains largely unchanged.